# Injecting Syntactic Sub-Networks in Transformer Language Models with Tree Regularization [Uncaptioned image]

### v1

@@ 9 Conclusion
We propose TreeReg, a structured regularizer that injects soft syntactic inductive biases into given circuits of a transformer LM, converting constituency parses of input sentences into differentiable orthogonality constraints on vector hidden states. Without requiring any architectural changes to the transformer, models pre-trained with TreeReg achieve up to 10% lower perplexities on out-of-distribution data and enhance syntactic generalization by up to 9.5 points on standard test suites. When applied to pre-trained LLMs, TreeReg improves syntactic generalization during continued pre-training and mitigates degradation on adversarial NLI benchmarks by up to 41.2 points when used during fine-tuning. TreeReg more than doubles the sample efficiency of syntactic generalization and does not require the entire training dataset to be parsed, making it practical and efficient. To the best of our knowledge, this work is among the first to translate insights from a mechanistic interpretability-related work (Murty et al., 2023c) into an actionable modeling innovation. We leave unsupervised alternatives to TreeReg, and pre-training at scale with TreeReg as future work.
@@
TreeReg requires constituency-parsed datasets, which might be difficult to obtain for languages other English due to the lack of easily available constituency parsers. This also makes TreeReg inapplicable to languages that do not have constituency structure. TreeReg also introduces some additional hyperparameters that require tuning, such as the layer and subset of attention heads on which TreeReg is be applied. As a heuristic, applying TreeReg at the middle layer of the model, on 25% of the attention heads, tends to work well across settings (see our ablations in Appendix F for more details). Finally, while we optimize TreeReg computation through vectorization, it still adds some computational overhead during training.
@@
Appendix A Algorithms for Computing TreeReg Loss and Induced Parse Trees
@@
function loss$(SCIS,P,i,j)$;
// Start: loss$(SCIS,P,0,n)$
@@
$SCIS$ : Dictionary of $\mathrm{SCIN}{}$ for all spans
@@
scores = $[SCIS(i,q)+SCIS(q+1,j)$ for $q\in[i,j-1]]$
@@
left_loss = loss$(SCIS,P,i,p)$
right_loss = loss$(SCIS,P,p+1,j)$
@@
function parse$(SCIS,i,j)$;
// Start: parse$(SCIS,0,n)$
@@
$SCIS$ : Dictionary of $\mathrm{SCIN}{}$ for all spans,
@@
split_point = $argmax_{q\in[i,j-1]}$ $SCIS(i,q)+SCIS(q+1,j)$
left_constituents = parse$(SCIS,i,split\_point)$
right_constituents = parse$(SCIS,split\_point+1,j)$
@@
Appendix B Efficient Computation of $\mathrm{SCIN}{}$
@@
Appendix C Dataset Statistics
@@
C.1 Tense Inflection Examples
@@
C.2 Question Formation Examples
@@ Appendix D NLI Setup
Appendix E Experiment Hyperparameters
@@
Appendix F Selection of Layer and Number of Attention Heads for TreeReg
@@
Appendix G Training on Randomized Parses
To assess if the gains from TreeReg are due to hierarchical computation guided by constituency parses or other factors, we conduct a controlled experiment. We train a 16-layer LM using the exact setup described in § 5.1, but instead of using gold parses, we provide TreeReg with random parses of BLLIP-LG sentences (Randomized TreeReg LM). These randomized parses are generated top-down recursively, by choosing random split points for each sentence span. We then compare performance with the TreeReg LM and Base LM on the BLiMP and SG test suites (Table 8).
Model BLiMP ($\uparrow$) SG ($\uparrow$)
Base LM 72.2 71.9
TreeReg LM 74.8 80
Randomized TreeReg LM 73.4 71.8
Table 8: Results on BLiMP and SG test suites for LMs trained from scratch on BLLIP-LG, with gold and randomized parses (Randomized TreeReg LM) passed to TreeReg. Performance decreases with randomized parses compared to gold parses, but improvements are still observed on BLiMP over the baseline LM.
As expected, the overall performance on the SG test suites drops marginally compared to the baseline when randomized parses are fed to TreeReg. On the other hand, BLiMP accuracies increase by 1.2 pts even in this setting. We hypothesize that this discrepancy arises from BLiMP’s less granular sentence-level perplexity-based approach, in contrast to the surprisal-based approach used in SG. This suggests that the SG test suites may be a more precise measure of syntactic generalization.

### v2

@@ 9 Conclusion
We propose TreeReg, a structured regularizer that injects soft syntactic inductive biases into given circuits of a transformer LM, converting constituency parses of input sentences into differentiable orthogonality constraints on vector hidden states. Without requiring any architectural changes to the transformer, models pre-trained with TreeReg achieve up to 10% lower perplexities on out-of-distribution data and enhance syntactic generalization by up to 9.5 points on standard test suites. When applied to pre-trained LLMs, TreeReg improves syntactic generalization during continued pre-training and mitigates degradation on adversarial NLI benchmarks by up to 41.2 points when used during fine-tuning. TreeReg more than doubles the sample efficiency of syntactic generalization and does not require the entire training dataset to be parsed, making it practical and efficient. To the best of our knowledge, this work is among the first to translate insights from a mechanistic interpretability-related work (Murty et al., 2023c) into an actionable modeling innovation. We leave unsupervised alternatives to TreeReg, pre-training at scale with TreeReg, and application of TreeReg to languages other than English as future work.
@@
TreeReg requires constituency-parsed datasets, which might be difficult to obtain for languages other English due to the lack of easily available constituency parsers. This also makes TreeReg inapplicable to languages that do not have constituency structure, and we limit the experiments presented here to English. TreeReg also introduces some additional hyperparameters that require tuning, such as the layer and subset of attention heads on which TreeReg is be applied. As a heuristic, applying TreeReg at the middle layer of the model, on 25% of the attention heads, tends to work well across settings (see our ablations in Appendix G for more details). Finally, while we optimize TreeReg computation through vectorization, it still adds some computational overhead during training.
Appendix A Discussion of Design Decisions
A.1 Last Hidden State as Representation for Prefix
We use the hidden state for the last token in a span as an easily computable representation of the span. There is some evidence suggesting that in the middle layers of an autoregressive transformer language model, the burden of computation of representations tends towards the right, indicating that intermediate hidden states summarize corresponding prefixes to some extent. For instance, Allen-Zhu and Li, 2023 find that when a transformer is trained on generations from a Probabilistic Context-Free Grammar (PCFG), ancestors of non-terminals are encoded at the hidden state of the last token in the span produced from the non-terminal. Therefore, we leverage the prefix information summarized in the hidden state of the last token of a span and rely on the LM’s substantial expressive capabilities to embed TreeReg’s orthogonality constraints into its hidden states.
A.2 Orthogonality for Operationalizing Contextual Independence
We choose to operationalize contextual independence between spans as orthogonality of correspoding hidden states, to design an efficiently computable loss function. There exist other ways of operationalizing contextual independence but these are inefficient and often require multiple forward passes on the transformer. For instance, Murty et al., 2023c defines contextual independence as the similarity between the sum of hidden states for a span with unconstrained attention, compared to those from a transformer with masked attention for the prefix of the span. Another efficient alternative is using the negative absolute value of cosine similarity between span vector representations to represent contextual independence. This does not work as well, resulting in a syntactic generalization of $76.4$ on SG test suites (a drop of $3.6$ points) and $70.6$ on BLiMP (a drop of $4.2$ points) when used during the pretraining of a 16-layer LM from scratch on BLLIP-LG (§ 5.1).
TreeReg softly encourages higher $\mathrm{SCIN}{}$ for constituents of the parse tree, and lower $\mathrm{SCIN}{}$ for other constituents. However, a low $\mathrm{SCIN}{}$ score indicates greater similarity between span vectors, due to a smaller orthogonal component between them, which can result in degenerate hidden states. In practice, this is mitigated by opposing pressure from $\mathcal{L}_{\text{LM}}{},$ that encourages divergent hidden states for different tokens in the sentence. To confirm this, we compute the average $\mathrm{SCIN}{}$ for constituent and non-constituent spans in the PennTreeBank test dataset, for the Base LM and TreeReg LM trained in § 5.1. For the TreeReg LM, the average $\mathrm{SCIN}{}$ for constituents is $1.29$, and that for non-constituents is $0.98$. In contrast, for the Base LM, the average $\mathrm{SCIN}{}$ for constituents is $0.66$, and that for non-constituents is $0.65$. We therefore conclude that the average $\mathrm{SCIN}{}$ for non-constituents is substantially positive for a model trained with TreeReg, and in fact is even higher than that for a model trained without TreeReg.
A.3 Computation of $\mathrm{SCIN}{}$
Our method for computation of $\mathrm{SCIN}{}$ involves the addition of two terms, the first one being $\|\mathrm{orth}(\bm{h}_{j},\bm{h}_{i-1})\|$, which captures the contextual independence between the information in a span and its prefix. The second term is $\|\mathrm{orth}(\bm{h}_{j+1},\bm{h}_{j})\|$, which is a measure of how much the suffix of a span is dependent on the information contained in the span. The information in constituent spans should be mostly self-contained in a tree-structured computation, and therefore the hidden state after the span is expected to be independent of it. While this could result in enforced orthogonalities between consecutive words that may belong to the same constituent, TreeReg only enforces soft constraints, and there is opposing pressure from $\mathcal{L}_{\text{LM}}$ mitigating any undesirable behavior. We also found that TreeReg improves syntactic generalization even without this future-masking constraint, reaching $77.4$ on SG test suites (a drop of $2.6$ points) and $72.8$ on BLiMP (a drop of $2.0$ points) when used during pre-training of a 16-layer LM from scratch on BLLIP-LG (§ 5.1).
A.4 Greedy Decoding for Recovering Parses
An alternative to the proposed top-down greedy decoding algorithm for recovering induced parse trees (§3.2) is a dynamic programming approach like the CKY algorithm (Cocke, 1969; Kasami, 1966; Younger, 1967). However, while computing $\mathcal{L}_{\text{TR}}{}$ during training, we only apply supervision on spans that are possible splits of constituents in the silver parse tree. For example, when applying TreeReg to “the quick brown fox jumped over the river”, there is no supervision on the $\mathrm{SCIN}{}$ for the span “fox jumped over”, as “fox” is a part of the constituent “the quick brown fox”, and “jumped over” is a part of the constituent “jumped over the river”. A globally optimal parsing algorithm such as CKY will also consider $\mathrm{SCIN}{}$ for spans that had no supervision on similar spans during training, and these can be arbitrarily large, resulting in worse parsing performance.
@@ A.4 Greedy Decoding for Recovering Parses
Appendix B Algorithms for Computing TreeReg Loss and Induced Parse Trees
@@
function loss$(SCIN,P,i,j)$;
// Start: loss$(SCIN,P,0,n)$
@@
$SCIN$ : Dictionary of $\mathrm{SCIN}{}$ for all spans
@@
scores = $[SCIN(i,q)+SCIN(q+1,j)$ for $q\in[i,j-1]]$
@@
left_loss = loss$(SCIN,P,i,p)$
right_loss = loss$(SCIN,P,p+1,j)$
@@
function parse$(SCIN,i,j)$;
// Start: parse$(SCIN,0,n)$
@@
$SCIN$ : Dictionary of $\mathrm{SCIN}{}$ for all spans,
@@
split_point = $argmax_{q\in[i,j-1]}$ $SCIN(i,q)+SCIN(q+1,j)$
left_constituents = parse$(SCIN,i,split\_point)$
right_constituents = parse$(SCIN,split\_point+1,j)$
@@
Appendix C Efficient Computation of $\mathrm{SCIN}{}$
@@
Appendix D Dataset Statistics
@@
D.1 Tense Inflection Examples
@@
D.2 Question Formation Examples
@@ Appendix E NLI Setup
Appendix F Experiment Hyperparameters
@@
Appendix G Selection of Layer and Number of Attention Heads for TreeReg